# Unsupervised Domain Adaptation for Binary Classification with an Unobservable Source Subpopulation

### v2

@@ 4.3 Model identification considerations
$+I_{0\cdot 1}{\rm log}\left\{\beta_{11}p_{11}({\bf x})+(1-\beta_{11}-{\boldsymbol{\beta}}^{\top}\mathbf{1})p_{01}({\bf x})\right\}+I_{0\cdot 0}{\rm log}\left\{{\boldsymbol{\beta}}^{\top}{\bf p}_{0}({\bf x})\right\}.$
In this formulation, the parameter with finite dimension is $\boldsymbol{\beta}$}. The model involves four nonparametric nuisance components: $p_{11}({\bf x})$, $p_{10}({\bf x})$, $p_{01}({\bf x})$, and $p_{00}({\bf x})$.
@@
The proof of Lemma 1 is provided in Appendix S.1. The identification conditions in Lemma 1 are intuitive and reasonable. If $\beta_{11}=0$, it degenerates to the situation that the source and target domains have the same support on both label $Y$ and background $A$, then the component $p_{11}({\bf x})$ is no longer relevant. Also, if $p_{10}({\bf x})=p_{00}({\bf x})$, the subpopulations of $(Y=1,A=0)$ and $(Y=0,A=0)$ become indistinguishable, and hence the individual probabilities $\beta_{10}$ and $\beta_{00}$ are not separately identifiable. Overall, these conditions are natural to ensure the problem is well-posed.
@@ 4.4 Estimating parameters of interest
To estimate the parameter $\boldsymbol{\beta}$}, we consider the distribution of attributes $\bf x$} in the subpopulation defined by $(R=0,A=0)$. By the law of total probability, we have
@@
Note that the distribution $p({\bf x}\mid R=0,A=0)$ is identifiable from the target population. The distributions $p_{10}({\bf x})$ and $p_{00}({\bf x})$ can be consistently estimated from the source population subgroups $(R=1,Y=1,A=0)$ and $(R=1,Y=0,A=0)$, respectively. Thus, the parameters ${\boldsymbol{\beta}}=(\beta_{00},\beta_{10})^{\rm T}$ can be estimated by minimizing a suitable discrepancy measure between the two sides of (9), such as an $L_{2}$ norm or a divergence-based criterion (e.g., Kullback–Leibler divergence), subject to the constraint in (10). Therefore, we reformulate the estimation of $\boldsymbol{\beta}$} as a constrained distribution matching problem:
$\widehat{\boldsymbol{\beta}}=\mbox{argmin}_{{\boldsymbol{\beta}}}D\left\{\widehat{p}({\bf x}\mid R=0,A=0)\middle\|\{\widehat{p}_{10}({\bf x})\beta_{10}+\widehat{p}_{00}({\bf x})\beta_{00}\}/\widehat{\hbox{pr}}(A=0|R=0)\right\},$ (11)
@@
The proof of Lemma 2 is provided in Appendix S.1. A key advantage of minimizing the KL divergence is that it circumvents the need to explicitly estimate the generative models $p_{10}({\bf x})$ and $p_{00}({\bf x})$, which are often difficult to model accurately in high dimensions. Instead, it suffices to estimate one background-specific prediction model $\xi_{0}({\bf x})$ using standard classification techniques on the source domain restricted to $A=0$. Finally, based on all of the above discussions, we summarize the implementation details of our proposed method in Algorithm 1.
@@
2:Estimated benchmark prediction models $\widehat{\xi}({\bf x})$, $\widehat{\xi}_{1}({\bf x})$ and $\widehat{\xi}_{0}({\bf x})$, proposed prediction models for the target $\widehat{\eta}({\bf x})$, $\widehat{\eta}_{1}({\bf x})$ and $\widehat{\eta}_{0}({\bf x})$; and subpopulation proportions $\widehat{\alpha}_{ya}$, $\widehat{\beta}_{ya}$.
@@
5:Estimate $\tau_{r}({\bf x})$ (defined in (6)) using data $\{({\bf X}_{i},A_{i},R_{i}=r):i=1,\cdots n_{r}\},r=0,1$, as $\widehat{\tau}_{r}({\bf x})$;
6:Estimate $\xi_{1}({\bf x})$ following (5), as $\widehat{\xi}_{1}({\bf x})$;
7:Estimate $\kappa({\bf x})$ (defined in (7)) using data $\{({\bf X}_{i},R_{i},A_{i}=1):i=1,\cdots n\}$, as $\widehat{\kappa}({\bf x})$;
8:Estimate $\boldsymbol{\beta}$} and $\alpha_{y,a}$ following (12) and (1), as $\widehat{\boldsymbol{\beta}}$ and $\widehat{\alpha}_{ya}$ for $(y,a)\in\{0,1\}$;
9:Estimate $\eta_{1}({\bf x})$, $\eta_{0}({\bf x})$ and $\eta({\bf x})$ following (8), as $\widehat{\eta}_{1}({\bf x})$, $\widehat{\eta}_{0}({\bf x})$ and $\widehat{\eta}({\bf x})$.
The above method adopts the idea of distribution matching. Alternatively, one may consider matching only certain moments rather than the full distribution. Due to space constraints, we defer the details to Appendix S.1.
4.5 Downstream tasks
With any loss function $\ell(\cdot)$, for the background-specific prediction model with $A=0$, the conditional risk is
where, for simplicity, we write $w(y)=\frac{\hbox{pr}(y\mid R=0,A=0)}{\hbox{pr}(y\mid R=1,A=0)}$. One can derive that $w(1)=\frac{\beta_{10}(\alpha_{00}+\alpha_{10})}{\alpha_{10}(\beta_{00}+\beta_{10})}$ and $w(0)=\frac{\beta_{00}(\alpha_{00}+\alpha_{10})}{\alpha_{00}(\beta_{00}+\beta_{10})}$. To evaluate the performance of the prediction model, it can be approximated as $\widehat{E}[\ell\{h({\bf X}),Y\}\,\widehat{w}(Y)\mid R=1,A=0]$. Furthermore, the model can be fine-tuned specifically for the target subgroup by minimizing the reweighted empirical risk:
where $\mathcal{F}$ is a suitable function class.
For the interest of space, for the other two prediction models, we only present the relations analogous to (13) without elaborations. For the background-specific prediction model with $A=1$, one can derive $E[\ell\{h({\bf X}),Y\}|R=0,A=1]$ as
For the overall prediction model, the conditional risk $E[\ell\{h({\bf X}),Y\}|R=0]$ is
@@ 5 Theoretical Results
For the interest of space, we only present the results for the background-specific prediction model with $A=0$. The results for the other two prediction models are parallel and can be similarly developed. To facilitate the analysis, we begin by formally defining the population-level (expected) objective function:
@@
Define ${\boldsymbol{f}}({\bf x})=\{f_{0}({\bf x}),f_{1}({\bf x})\}^{\rm T}$, where $f_{0}({\bf x})={\rm log}\{\xi_{0}({\bf x})\}-\frac{1}{2}[{\rm log}\{\xi_{0}({\bf x})\}+{\rm log}\{1-\xi_{0}({\bf x})\}]$ and $f_{1}({\bf x})={\rm log}\{1-\xi_{0}({\bf x})\}-\frac{1}{2}[{\rm log}\{\xi_{0}({\bf x})\}+{\rm log}\{1-\xi_{0}({\bf x})\}]$, and the corresponding estimate is $\{\widehat{f}_{k}({\bf x})\}_{k=0}^{1}$. There exist a constant $c>0$ and a sequence $r_{n_{1\cdot 0}}\to 0$ such that, for almost every $\bf x$}, we have
$\hbox{pr}\left(\|\widehat{{\boldsymbol{f}}}({\bf x})-{\boldsymbol{f}}({\bf x})\|_{2}>t\right)\leq\exp\left\{-t^{2}/(c^{2}r_{n_{1\cdot 0}}^{2})\right\},\quad\forall\ t>0.$
@@
$\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{1}\leq c_{10}\chi_{n}\sqrt{{\rm log}(1/\delta)}.$
The proof of Theorem 1 is provided in Appendix S.2. Theorem 1 establishes the consistency of the estimator $\widehat{{\boldsymbol{\beta}}}$, provided that $r_{n_{1\cdot 0}}\sqrt{{\rm log}(n_{0\cdot 0})}\to 0$ as $n_{1\cdot 0},n_{0\cdot 0}\to\infty$.
Next, we establish a generalization bound for the fitted model (14), which is obtained via weighted empirical risk minimization over the source subgroup. Let $\mathcal{F}$ denote the hypothesis class of classifiers. For any $h\in\mathcal{F}$ and a weight function $w(y):y\rightarrow{\bf R}$, we define the population-level weighted loss and its empirical counterpart based on the source subgroup data as follows:
@@
To establish our generalization bound, we utilize the concept of Rademacher complexity (Bartlett and Mendelson, 2002), denoted as $\mathcal{R}_{n}(\mathcal{G})$ (see Appendix S.2 for details), and impose the following assumption on the loss function:
@@
We now present the generalization bound for the learned model, with its proof provided in Appendix S.2.
@@
${\mathcal{L}}_{0}(\widehat{h}_{\widehat{w}})-\min_{h\in\mathcal{F}}{\mathcal{L}}_{0}(h)\leq 2\mathcal{R}_{n_{1\cdot 0}}(\mathcal{G})+dB\|\widehat{\boldsymbol{\beta}}-{\boldsymbol{\beta}}\|_{1}+c\left\{\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{0}}}\right\},$
where $\mathcal{G}=\{w(y)\ell\{h({\bf x}),y\}:h\in\mathcal{F}\}$, and $\mathcal{R}_{n_{1\cdot 0}}(\mathcal{G})$ denotes its Rademacher complexity as defined in Appendix S.2.
@@
Proposition 2 indicates that the generalization bound depends on the estimation error $\|\widehat{\boldsymbol{\beta}}-{\boldsymbol{\beta}}\|_{1}$, which can be directly controlled based on the conditions listed in Assumption 1, implying that different estimation procedures for $\boldsymbol{\beta}$} will yield different upper bounds. In Theorem 1, we established an upper bound for the estimation error of $\widehat{{\boldsymbol{\beta}}}$, which directly leads to a refined generalization bound for the learned classifier $\widehat{h}_{\widehat{w}}$. Specifically, for any $\delta>0$, with probability at least $1-13\delta$, the following inequality holds:
@@
where $c_{10}$ is the constant appearing in Theorem 1, and $\alpha_{n}$ characterizes the convergence rate of $\widehat{{\boldsymbol{\beta}}}$.
@@ 6 Synthetic Data Results
We consider a structured data-generating process in which the covariates ${\bf X}\in{\bf R}^{4}$ are drawn from a distribution conditioned on a latent pair $(Y,A)$, where $Y\in\{0,1\}$ denotes the class label and $A\in\{0,1\}$ denotes the background. The generation begins by sampling $(Y,A)$ according to a predefined distribution.
@@
In the source domain, we consider $(Y,A)\in\{(0,0),(0,1),(1,0)\}$, each occurring with probability $1/3$. The covariates ${\bf X}\in{\bf R}^{4}$ are generated as ${\bf X}\sim N(\boldsymbol{\mu}_{\rm{YA}},{\bf I}_{4})$, where $\boldsymbol{\mu}_{\rm{YA}}$ denotes the mean vector for each combination and ${\bf I}_{4}$ is the $4\times 4$ identity matrix. The stratum $(1,1)$ is excluded from the source. In the target domain, all four combinations $(Y,A)\in\{0,1\}^{2}$ appear with equal probability $1/4$, and $\bf X$} is drawn from the same distribution $N(\boldsymbol{\mu}_{\rm{YA}},{\bf I}_{4})$ with distinct means:
$\boldsymbol{\mu}_{00}=(1,0,0,0)^{\rm T},\quad\boldsymbol{\mu}_{01}=(0,0,1,0)^{\rm T},\quad\boldsymbol{\mu}_{10}=(0,1,0,0)^{\rm T},\quad\boldsymbol{\mu}_{11}=(0,0,0,1)^{\rm T}.$
For model estimation, we utilize the known data-generating distribution to compute the conditional probability models required by both our proposed estimators and the benchmark method. Specifically, we calculate the five key conditional probabilities needed for implementation: $\xi_{0}({\bf x})$, $\xi({\bf x})$, $\tau_{0}({\bf x})$, $\tau_{1}({\bf x})$, and $\kappa({\bf x})$, which together determine the models $\eta_{0}({\bf x})$, $\eta_{1}({\bf x})$ and $\eta({\bf x})$ for our method, and $\xi_{0}({\bf x})$, $\xi_{1}({\bf x})$ and $\xi({\bf x})$ for the benchmark. Additionally, the parameters $\{\beta_{ya}:y=0,1;a=0,1\}$ and $\{\alpha_{ya}:y=0,1;a=0,1\}$ are treated as known. The classification threshold is set to $0.5$, consistent with the standard Bayesian decision rule; that is, a sample is classified as positive if the predicted probability exceeds $0.5$.
@@
To assess the performance of the proposed estimators, we conduct 100 simulations for each configuration and report the results using boxplots that compare $\{\widehat{\eta}_{1}({\bf x}),\widehat{\xi}_{1}({\bf x})\}$ and $\{\widehat{\eta}({\bf x}),\widehat{\xi}({\bf x})\}$ across varying sample sizes. The left panel of Figure 1 illustrates the performance of $\{\widehat{\eta}_{1}({\bf x}),\widehat{\xi}_{1}({\bf x})\}$ for $n_{0}=1000$ and $6000$, with $n_{1}$ ranging from $1000$ to $8000$. The right panel of Figure 1 shows the corresponding performance of $\{\widehat{\eta}({\bf x}),\widehat{\xi}({\bf x})\}$ under the same settings. Performance is evaluated using two standard metrics: accuracy and F1 score. In both cases, the proposed estimators consistently outperform the benchmark estimators. Moreover, as $n_{0}$ increases, the variance of the estimators decreases, suggesting improved stability and reliability. In particular, for the F1 score of $\widehat{\eta}({\bf x})$, performance steadily improves as $n_{0}$ increases, further demonstrating the robustness and effectiveness of the proposed method. Comparable patterns are also observed when $n_{1}$ is fixed at $1000$ and $6000$ while $n_{0}$ varies from $1000$ to $8000$, as shown in Figure 2.
Figure 1: The left panel displays the performance of the F1 score and accuracy for $\eta_{1}({\bf x})$ and $\xi_{1}({\bf x})$ across different scenarios, while the right panel presents the corresponding results for $\eta({\bf x})$ and $\xi({\bf x})$.
Figure 2: The left panel displays the performance of the F1 score and accuracy for $\eta_{1}({\bf x})$ and $\xi_{1}({\bf x})$ across different scenarios, while the right panel presents the corresponding results for $\eta({\bf x})$ and $\xi({\bf x})$.

### v3

@@ 4.3 Model identification considerations
$+I_{0\cdot 1}{\rm log}\left\{\beta_{11}p_{11}({\bf x})+(1-\beta_{11}-{\bm{\beta}}^{\top}\mathbf{1})p_{01}({\bf x})\right\}+I_{0\cdot 0}{\rm log}\left\{{\bm{\beta}}^{\top}{\bf p}_{0}({\bf x})\right\}.$
In this formulation, the parameter with finite dimension is $\bm{\beta}$}. The model involves four nonparametric nuisance components: $p_{11}({\bf x})$, $p_{10}({\bf x})$, $p_{01}({\bf x})$, and $p_{00}({\bf x})$.
@@
The proof of Lemma 1 is provided in Appendix A.1. The identification conditions in Lemma 1 are intuitive and reasonable. If $\beta_{11}=0$, it degenerates to the situation that the source and target domains have the same support on both label $Y$ and background $A$, then the component $p_{11}({\bf x})$ is no longer relevant. Also, if $p_{10}({\bf x})=p_{00}({\bf x})$, the subpopulations of $(Y=1,A=0)$ and $(Y=0,A=0)$ become indistinguishable, and hence the individual probabilities $\beta_{10}$ and $\beta_{00}$ are not separately identifiable. Overall, these conditions are natural to ensure the problem is well-posed.
@@ 4.4 Estimating parameters of interest
To estimate the parameter $\bm{\beta}$}, we consider the distribution of attributes $\bf x$} in the subpopulation defined by $(R=0,A=0)$. By the law of total probability, we have
@@
Note that the distribution $p({\bf x}\mid R=0,A=0)$ is identifiable from the target population. The distributions $p_{10}({\bf x})$ and $p_{00}({\bf x})$ can be consistently estimated from the source population subgroups $(R=1,Y=1,A=0)$ and $(R=1,Y=0,A=0)$, respectively. Thus, the parameters ${\bm{\beta}}=(\beta_{00},\beta_{10})^{\rm T}$ can be estimated by minimizing a suitable discrepancy measure between the two sides of (8), such as an $L_{2}$ norm or a divergence-based criterion (e.g., Kullback–Leibler divergence), subject to the constraint in (9). Therefore, we reformulate the estimation of $\bm{\beta}$} as a constrained distribution matching problem:
$\widehat{\bm{\beta}}=\mbox{argmin}_{{\bm{\beta}}}D\left\{\widehat{p}({\bf x}\mid R=0,A=0)\middle\|\{\widehat{p}_{10}({\bf x})\beta_{10}+\widehat{p}_{00}({\bf x})\beta_{00}\}/\widehat{\hbox{pr}}(A=0|R=0)\right\},$ (10)
@@
The proof of Lemma 2 is provided in Appendix A.1. A key advantage of minimizing the KL divergence is that it circumvents the need to explicitly estimate the generative probabilities $p_{10}({\bf x})$ and $p_{00}({\bf x})$, which are often difficult to model accurately in high dimensions. Instead, it suffices to estimate one background-specific prediction probability $\xi_{0}({\bf x})$ using standard classification techniques on the source domain restricted to $A=0$.
Finally, based on all of the above discussions, we summarize the implementation details of our proposed method in Algorithm 1.
@@
2:Estimated benchmark predictive probabilities $\widehat{\xi}({\bf x})$ and $\widehat{\xi}_{0}({\bf x})$, proposed predictive probabilities for the target $\widehat{\eta}({\bf x})$, $\widehat{\eta}_{1}({\bf x})$ and $\widehat{\eta}_{0}({\bf x})$; and subpopulation proportions $\widehat{\alpha}_{ya}$, $\widehat{\beta}_{ya}$.
@@
5:Estimate $\tau_{r}({\bf x})$ (defined in Proposition 1) using data $\{({\bf X}_{i},A_{i},R_{i}=r):i=1,\cdots n_{r}\},r=0,1$, as $\widehat{\tau}_{r}({\bf x})$;
6:Estimate $\kappa({\bf x})$ (defined in (6)) using data $\{({\bf X}_{i},R_{i},A_{i}=1):i=1,\cdots n\}$, as $\widehat{\kappa}({\bf x})$;
7:Estimate $\bm{\beta}$} and $\alpha_{y,a}$ following (11) and (1), as $\widehat{\bm{\beta}}$ and $\widehat{\alpha}_{ya}$ for $(y,a)\in\{0,1\}$;
8:Estimate $\eta_{1}({\bf x})$, $\eta_{0}({\bf x})$ and $\eta({\bf x})$ following (7), as $\widehat{\eta}_{1}({\bf x})$, $\widehat{\eta}_{0}({\bf x})$ and $\widehat{\eta}({\bf x})$.
The above method adopts the idea of distribution matching. Alternatively, one may consider matching only certain moments rather than the full distribution. Due to space constraints, we defer the details to Appendix A.1.
@@ 5 Theoretical Results
For the interest of space, we only present the results for the background-specific predictive probability with $A=0$. The results for the other two predictive probabilities are parallel and can be similarly developed. To facilitate the analysis, we begin by formally defining the population-level (expected) objective function:
@@
Define ${\bm{f}}({\bf x})=\{f_{0}({\bf x}),f_{1}({\bf x})\}^{\rm T}$, where $f_{0}({\bf x})={\rm log}\{\xi_{0}({\bf x})\}-\frac{1}{2}[{\rm log}\{\xi_{0}({\bf x})\}+{\rm log}\{1-\xi_{0}({\bf x})\}]$ and $f_{1}({\bf x})={\rm log}\{1-\xi_{0}({\bf x})\}-\frac{1}{2}[{\rm log}\{\xi_{0}({\bf x})\}+{\rm log}\{1-\xi_{0}({\bf x})\}]$, and the corresponding estimate is $\{\widehat{f}_{k}({\bf x})\}_{k=0}^{1}$. There exist a constant $c>0$ and a sequence $r_{n_{1\cdot 0}}\to 0$ such that, for almost every $\bf x$}, we have
$\hbox{pr}\left(\|\widehat{{\bm{f}}}({\bf x})-{\bm{f}}({\bf x})\|_{2}>t\right)\leq\exp\left\{-t^{2}/(c^{2}r_{n_{1\cdot 0}}^{2})\right\},\quad\forall\ t>0.$
@@
$\|\widehat{\bm{\beta}}-\bm{\beta}\|_{1}\leq c_{10}\chi_{n}\sqrt{{\rm log}(1/\delta)}.$
The proof of Theorem 1 is provided in Appendix A.2. Theorem 1 establishes the consistency of the estimator $\widehat{{\bm{\beta}}}$, provided that $r_{n_{1\cdot 0}}\sqrt{{\rm log}(n_{0\cdot 0})}\to 0$ as $n_{1\cdot 0},n_{0\cdot 0}\to\infty$.
Theorem 1 establishes the upper bound of the parameter estimate. Beyond parameter estimation, an important question is how the resulting estimator performs in downstream prediction tasks.
With any loss function $\ell(\cdot)$ and some function $h(\cdot)$, for the background-specific prediction model with $A=0$, the conditional risk is
where, for simplicity, we write $w(y)=\frac{\hbox{pr}(y\mid R=0,A=0)}{\hbox{pr}(y\mid R=1,A=0)}$. One can derive that $w(1)=\frac{\beta_{10}(\alpha_{00}+\alpha_{10})}{\alpha_{10}(\beta_{00}+\beta_{10})}$ and $w(0)=\frac{\beta_{00}(\alpha_{00}+\alpha_{10})}{\alpha_{00}(\beta_{00}+\beta_{10})}$. To evaluate the performance of the predictive probability, it can be approximated as $\widehat{E}[\ell\{h({\bf X}),Y\}\,\widehat{w}(Y)\mid R=1,A=0]$. Furthermore, the model can be fine-tuned specifically for the target subgroup by minimizing the reweighted empirical risk:
where $\mathcal{F}$ is a suitable function class. For the other two predictive probabilities, the relations analogous to (12) can be derived similarly. For the background-specific predictive probability with $A=1$, one can derive $E[\ell\{h({\bf X}),Y\}|R=0,A=1]$ as
For the overall predictive probability, the conditional risk $E[\ell\{h({\bf X}),Y\}|R=0]$ is
The above relations show that the target-domain risk can be expressed in terms of appropriately reweighted risks computed from the observed source data. This motivates the use of weighted empirical risk minimization for learning classifiers tailored to the target domain.
We now establish a generalization bound for the fitted model (13), which is obtained via weighted empirical risk minimization over the source subgroup. Let $\mathcal{F}$ denote the hypothesis class of classifiers. For any $h\in\mathcal{F}$ and a weight function $w(y):y\rightarrow{\bf R}$, we define the population-level weighted loss and its empirical counterpart based on the source subgroup data as follows:
@@
To establish our generalization bound, we utilize the concept of Rademacher complexity (Bartlett and Mendelson, 2002), denoted as $\mathcal{R}_{n}(\mathcal{G})$ (see Appendix A.2 for details), and impose the following assumption on the loss function:
@@
We now present the generalization bound for the learned model, with its proof provided in Appendix A.2.
@@
${\mathcal{L}}_{0}(\widehat{h}_{\widehat{w}})-\min_{h\in\mathcal{F}}{\mathcal{L}}_{0}(h)\leq 2\mathcal{R}_{n_{1\cdot 0}}(\mathcal{G})+dB\|\widehat{\bm{\beta}}-{\bm{\beta}}\|_{1}+c\left\{\sqrt{\frac{{\rm log}(1/\delta)}{n_{1\cdot 0}}}+\sqrt{\frac{{\rm log}(1/\delta)}{n_{0}}}\right\},$
where $\mathcal{G}=\{w(y)\ell\{h({\bf x}),y\}:h\in\mathcal{F}\}$, and $\mathcal{R}_{n_{1\cdot 0}}(\mathcal{G})$ denotes its Rademacher complexity as defined in Appendix A.2.
@@
Proposition 2 indicates that the generalization bound depends on the estimation error $\|\widehat{\bm{\beta}}-{\bm{\beta}}\|_{1}$, which can be directly controlled based on the conditions listed in Assumption 1, implying that different estimation procedures for $\bm{\beta}$} will yield different upper bounds. In Theorem 1, we established an upper bound for the estimation error of $\widehat{{\bm{\beta}}}$, which directly leads to a refined generalization bound for the learned classifier $\widehat{h}_{\widehat{w}}$. Specifically, for any $\delta>0$, with probability at least $1-13\delta$, the following inequality holds:
@@
where $c_{10}$ is the constant appearing in Theorem 1, and $\alpha_{n}$ characterizes the convergence rate of $\widehat{{\bm{\beta}}}$.
@@ 6 Synthetic Data Results
We consider a structured data-generating process in which the covariate ${\bf X}\in{\bf R}^{4}$ is drawn from a distribution conditioned on a pair $(Y,A)$, where $Y\in\{0,1\}$ denotes the class label and $A\in\{0,1\}$ denotes the background. The generation begins by sampling $(Y,A)$ according to a predefined distribution.
@@
In the source domain, we consider $(Y,A)\in\{(0,0),(0,1),(1,0)\}$, each occurring with probability $1/3$. The covariate ${\bf X}\in{\bf R}^{4}$ is generated as ${\bf X}\sim N(\bm{\mu}_{\rm{YA}},{\bf I}_{4})$, where $\bm{\mu}_{\rm{YA}}$ denotes the mean vector for each combination and ${\bf I}_{4}$ is the $4\times 4$ identity matrix. The stratum $(1,1)$ is excluded from the source. In the target domain, all four combinations $(Y,A)\in\{0,1\}^{2}$ appear with equal probability $1/4$, and $\bf X$} is drawn from the same distribution $N(\bm{\mu}_{\rm{YA}},{\bf I}_{4})$ with distinct means:
$\bm{\mu}_{00}=(1,0,0,0)^{\rm T},\quad\bm{\mu}_{01}=(0,0,1,0)^{\rm T},\quad\bm{\mu}_{10}=(0,1,0,0)^{\rm T},\quad\bm{\mu}_{11}=(0,0,0,1)^{\rm T}.$
We follow Algorithm 1 to compute the conditional probabilities required by the proposed approach as well as the two naive benchmarks. Specifically, we calculate the five key conditional probabilities needed for implementation: $\xi_{0}({\bf x})$, $\xi({\bf x})$, $\tau_{0}({\bf x})$, $\tau_{1}({\bf x})$, and $\kappa({\bf x})$, which, together with the estimators of the parameters $\{\beta_{ya}:y=0,1;a=0,1\}$ and $\{\alpha_{ya}:y=0,1;a=0,1\}$, determine the predictive probabilities $\eta_{0}({\bf x})$, $\eta_{1}({\bf x})$ and $\eta({\bf x})$ for the proposed method. In addition, we compute the naive benchmark $\gamma({\bf x})$ following (5), which corresponds to the standard UDA method relying solely on the label shift assumption without accounting for the subgroup information.
@@
To assess the performance of the proposed approach as well as to compare with the two naive benchmarks, we conduct 100 simulations for each configuration and summarize the results using boxplots that compare $\widehat{\eta}({\bf x}),\widehat{\xi}({\bf x}),\widehat{\gamma}({\bf x})$} across varying sample sizes. The left panel of Figure 1 displays the performance of $\widehat{\eta}({\bf x}),\widehat{\xi}({\bf x})$ and $\widehat{\gamma}({\bf x})$ for $n_{0}=1000$ and $6000$, with $n_{1}$ ranging from $1000$ to $8000$, while the right panel shows the corresponding results for $n_{1}=1000$ and $6000$, with $n_{0}$ ranging from $1000$ to $8000$. Performance is evaluated using two standard metrics: accuracy and F1 score. In both metrics, the proposed approach consistently outperforms the benchmark estimators. Notably, the accuracy of $\widehat{\eta}({\bf x})$ steadily improves with larger $n_{0}$ (or $n_{1}$), further demonstrating the robustness and effectiveness of the proposed method.
Figure 1: Left panel: Accuracy and F1 score of $\widehat{\eta}({\bf x})$, $\widehat{\xi}({\bf x})$, and $\widehat{\gamma}({\bf x})$ with fixed $n_{0}=1000/6000$ but $n_{1}$ varies. Right panel: Accuracy and F1 score of $\widehat{\eta}({\bf x})$, $\widehat{\xi}({\bf x})$, and $\widehat{\gamma}({\bf x})$ with fixed $n_{1}=1000/6000$ but $n_{0}$ varies.